OpenStudy (anonymous):

salt poured from conveyor belt at 30cft/min. foms a conical pile /w a circular basewhose ht. and base diam. are equal. how fast is ht. of pile incr. when ht. is 10 ft.

OpenStudy (dumbcow):

Given: h = 2r , V = pi*r^2*h since we want rate the height is changing, get volume in terms of height \[V = \pi (\frac{h}{2})^{2} h = \frac{\pi h^{3}}{4}\] Now set up a related rates equation \[\frac{dh}{dt} = \frac{dV}{dt}*\frac{dh}{dV}\] \[\frac{dV}{dt} = 30\] take derivative of V wrt h \[\frac{dV}{dh} = \frac{3 \pi h^{2}}{4} \rightarrow \frac{dh}{dV} = \frac{4}{3 \pi h^{2}}\] thus \[\frac{dh}{dt} = \frac{40}{\pi h^{2}}\]